OpenStudy (anonymous):

is (8,-6) a solution to 8x-6y=7

OpenStudy (anonymous):

Substitue x=8 and y=-6 in that equation and see if it holds.

OpenStudy (anonymous):

or graph the line and see if that point exists on the line </more work than needed>

OpenStudy (anonymous):

I worked 8(8) -6(-6) = 64-36= 28 so the answer would be no, is that right?

Directrix (directrix):

8(8) - 6(-6) = 64 + 36 = 100 100 is not equal to 7. So, no. Note: -6(-6) = + 36 and not - 36